OpenStudy (anonymous):

Determine the solutions to the quadratic equation -4x2 - 32x - 64 = 0.

OpenStudy (anonymous):

140

mathslover (mathslover):

a= - 4 b = -32 c = -64 1st root : \[32 + \sqrt{1024+1024}/2(-4)\] \[32 + \sqrt{2(1024)}/-8\] \[32 + 32\sqrt{2}/-8\] \[32(1+\sqrt{2}/-8\] \[-4{1+\sqrt{2}}\] 2nd root : \[32 - \sqrt{1024+1024}/2(-4)\] \[32 -\sqrt{2(1024)}/-8\] \[32-32\sqrt{2}/-8\] \[-4(1-\sqrt{2})\]

OpenStudy (anonymous):

Thank you.